# A Data-Centric Approach To Generate Faithful and High Quality Patient Summaries with Large Language Models

### v1

@@ Abstract
Patients often face difficulties in understanding their hospitalizations, while healthcare workers have limited resources to provide explanations. In this work, we investigate the potential of large language models to generate patient summaries based on doctors’ notes and study the effect of training data on the faithfulness and quality of the generated summaries. To this end, we develop a rigorous labeling protocol for hallucinations, and have two medical experts annotate 100 real-world summaries and 100 generated summaries. We show that fine-tuning on hallucination-free data effectively reduces hallucinations from 2.60 to 1.55 per summary for Llama 2, while preserving relevant information. Although the effect is still present, it is much smaller for GPT-4 when prompted with five examples (0.70 to 0.40). We also conduct a qualitative evaluation using hallucination-free and improved training data. GPT-4 shows very good results even in the zero-shot setting. We find that common quantitative metrics do not correlate well with faithfulness and quality. Finally, we test GPT-4 for automatic hallucination detection, which yields promising results.
@@
We use the MIMIC-IV-Note dataset (Johnson et al., 2023; Goldberger et al., 2000). We create a filtered dataset of discharge instructions (MIMIC-IV-Note-DI) and annotate hallucinations in real-world and generated patient summaries (Hallucinations-{MIMIC,Generated}-DI). We will make all datasets available on PhysioNet. Our code is available on Github: https://github.com/stefanhgm/patient_summaries_with_llms
@@ 1 Introduction
Figure 1: Overview of this study. We created the filtered dataset MIMIC-IV-Note-DI and labeled hallucinations in 100 examples (original). We created a version without hallucinations (cleaned) and checked if training on this data reduced hallucination for Llama 70B and GPT-4. We also created a version with fixed mistakes (cleaned & improved) that we used for a qualitative evaluation. [^2] Llama icon from: vecteezy.com/free-vector/llama-icon
@@
Many patients do not understand the events that occurred during their hospitalization and the subsequent actions they need to take (Kebede et al., 2014). For instance, Horwitz et al. (2013) performed post-discharge interviews and found that only 59.6% of the patients were able to accurately describe their admission diagnosis and 43.9% could fully describe their scheduled follow-up appointments. An improved discharge communication is associated with lower hospital readmission rates and higher adherence to treatment regiment (Becker et al., 2021). A potential intervention to improve patient informedness could be patient-oriented summaries that describe all relevant facts in layperson language (Federman et al., 2018). However, writing high-quality patient summaries is a difficult and time-consuming task (Mueller et al., 2015) and healthcare workers already face high workloads (Phillips, 2020; Watson et al., 2019).
@@
Large language models (LLMs) have demonstrated strong capabilities on many natural language tasks including medical summarization (Van Veen et al., 2023b). However, LLMs are prone to generate unsupported or erroneous facts also referred to as hallucinations (Maynez et al., 2020). In healthcare, this issue is further aggravated by the fragmented nature of healthcare data as datasets often do not perfectly mimic the data available at the point of care. For example, datasets for medical summarization may not include the full patient history to accompany the written summarization leading to “hallucinations” in the human-written summary. Training or fine-tuning on this data replicates these artifacts. Several techniques for preventing hallucinations have been studied (Huang et al., 2023). However, hallucinations can highly vary in complexity escaping automatic detection and making careful human annotation necessary (Thomson et al., 2023). This also applies to medical summaries (Moramarco et al., 2022).
@@
Recent work has emphasized the sample efficiency of LLMs. Zhou et al. (2023) showed that 1,000 carefully selected fine-tuning examples sufficed for successful alignment of Llama 2 (Touvron et al., 2023). Also, GPT-4 achieved state-of-the-art performance on many benchmarks using at most 25 in-context examples (OpenAI et al., 2023). These findings open up opportunities for data-centric approaches that leverage data curated by human domain experts. In this study, we create 100 training examples without hallucinations and 100 training examples with improved quality to study their effect on the faithfulness and quality of LLMs. Our main contributions are:
  • We create the filtered dataset MIMIC-IV-Note-DI, which consists of 100,175 hospital courses and patient summaries written by doctors.
  • We introduce a labeling protocol for hallucinations and two medical experts annotate 100 real-world and 100 generated patient summaries.
@@
  • We demonstrate that fine-tuning on data with manually removed hallucinations can effectively reduce hallucinations for LLMs while preserving key information in patient summaries.
  • We conduct a qualitative evaluation that shows GPT-4 outperforms Llama 2 and often surpasses the quality of human-written summaries.
  • We introduce an automatic hallucination detection task based on our annotations and provide a specialized prompt and results for GPT-4.
@@ 2 Related Work
Given the large burden of and repetition within clinical documentation, automated clinical summarization is of great interest across healthcare; generation targets include discharge summaries, radiology impression sections, patient-facing instructions, and problem lists (Van Veen et al., 2023b; Adams et al., 2022). Consequently, there has been significant research in the past few years exploring how the advent of LLMs can be leveraged to automate clinical summarization. Approaches for control have included both zero- and few-shot prompting (Van Veen et al., 2023b) to fine-tuning (Adams et al., 2022, 2023; Moramarco et al., 2022; Cai et al., 2022; Van Veen et al., 2023a). Results have been promising, e.g., Van Veen et al. (2023b) finding that GPT-4 is often preferred over human-generated summaries in terms of correctness and completeness.
@@
However, hallucinations – the generation of facts not grounded in the original text — are a recurring concern in text summarization, both across natural language processing (NLP) in general and specifically in the clinical domain (Zhang et al., 2020b; Kryscinski et al., 2020; Adams et al., 2022; Cai et al., 2022; Xie et al., 2023). Given the ubiquity of this issue, several taxonomies have sprung up in the wider NLP community for producing fine-grained annotations of the accuracy of generated text (Thomson and Reiter, 2020; Mishra et al., 2024). In the clinical domain, there have been similar efforts to annotate hallucinations in generated text, but this work has often occurred on proprietary data, with coarse buckets, or without release of annotations (Adams et al., 2023; Cai et al., 2022; Van Veen et al., 2023b). The closest public release was a set of edits clinicians made to summaries auto-generated from transcripts, but this did not target hallucinations specifically (Moramarco et al., 2022). To our knowledge, this work is first to release a dataset highlighting hallucinations in clinical summarization. These are particularly important because automated metrics of summarization quality, while fast to compute, often do not correlate with manually evaluated performance, a finding also replicated in this work (Van Veen et al., 2023b).
@@
There have been three major strategies to mitigate the effect of hallucinations in auto-generated summaries. The first is to conduct post-hoc detection of hallucinations. This has largely involved learning a model of hallucinations from synthetically generated data (Cai et al., 2022; Zhou et al., 2021; Kryscinski et al., 2020) or identifying normalized concepts (e.g., UMLS concepts) that appear in the summary, but not the source (Nan et al., 2021; Adams et al., 2022). The second strategy is to minimize hallucinations by changing the underlying generation model, e.g., by grounding in Wikipedia, or first converting the source context as triples as an intermediary before generation (Zhang et al., 2020b; Tian et al., 2023; Semnani et al., 2023; Aralikatte et al., 2021; Cao et al., 2018; Cao and Wang, 2021). The final strategy is to improve the quality of the training data itself, either by removing poor training examples (Nan et al., 2021) or by improving the quality of the training examples. For example, to avoid decreasing sample size, Adams et al. (2022) revised the reference text via contrastive learning on synthetic data. In this work, we show that for sample efficient LLMs, it is sufficient to revise only a small number of training examples that are feasible even manually.
Quantity Value (SD)
MIMIC-IV-Note-DI (100,175 context-summary pairs)
Full context # Sentences $118.2$ ($50.4$)
# Words $2088.8$ ($778.1$)
# Tokens $4367.1$ ($1625.3$)
Short context (BHC) [used in this work] # Sentences $33.0$ ($19.0$)
# Words $552.0$ ($314.0$)
# Tokens $858.6$ ($498.3$)
Summaries (DI) # Sentences $6.5$ ($2.6$)
# Words $113.2$ ($47.4$)
# Tokens $145.4$ ($61.4$)
Table 1: Overview of MIMIC-IV-Note-DI. There are two versions: Full context contains all notes prior to the Discharge Instruction (DI). Short context only has the Brief Hospital Course (BHC). [^4] Sentences and words were determined with nltk. Tokens were determined with the Llama 2 tokenizer. Full version see Table 10.
@@ 3 Datasets
3.1 MIMIC-IV-Note-DI
We created a dataset of diverse real-world patient summaries paired with the corresponding doctor’s notes as context. MIMIC-IV-Note v2.2 includes 331,793 deidentified free-text clinical notes from 145,915 patients admitted to the Beth Israel Deaconess Medical Center in Boston, MA, USA (Johnson et al., 2023; Goldberger et al., 2000). Each note is composed of various sections that describe a patient’s hospital course. For this work, we selected the section Discharge Instructions (DI) as the summaries because it provides a concise summary for the patient. The section Brief Hospital Course (BHC) was used as context. It contains the most relevant information about the hospital course written for medical professionals (see Figure 2). We chose this shortened context to reduce the effort for the human annotators and to better fit it into the models’ context windows. Many DI contained irrelevant artifacts that could distort the downstream error analysis. Hence, we designed a preprocessing pipeline that filtered poor summaries and removed irrelevant content (see Appendix A). We release two versions of the MIMIC-IV-Note-DI (MIMIC) dataset: one with the full context and another with the BHC as context (see Table 4).
3.2 MIMIC Data for Hallucination Labeling
To annotate hallucinations in 100 MIMIC summaries, we considered a subset of the MIMIC-IV-Note-DI dataset. This subset was filtered to include contexts with a length of at most 4,000 characters and summaries with at least 600 characters. This was done to reduce the amount of context to take into account for the annotators and to increase the information in the summaries. The resulting subset contained 26,178 of the original 100,175 examples. Our quantitative evaluations were also performed on this subset to verify that the selection did not introduce a significant bias (see Table 12). From the subset of 26,178 examples, we randomly selected 10 examples to train the annotators and 100 examples for labeling hallucinations.
We derived two datasets from these 100 annotated MIMIC examples (Original). These two datasets were used for our data-centric hallucination reduction experiments and qualitative evaluation (see Figure 2):
  • Original: 100 examples randomly chosen from the subset of MIMIC-IV-Note-DI used for labeling hallucinations in MIMIC summaries.
@@ 3.2 MIMIC Data for Hallucination Labeling
  • Cleaned: 100 Original examples with labeled hallucinations manually removed or replaced.
  • Cleaned & Improved: 100 Cleaned examples with mistakes corrected and artifacts removed.
@@ 4.1 Generation of Patient Summaries
In this work, we generated patient summaries given the BHC as context. Formally, we have a set of contexts $\mathcal{C}=\{C_{1},...,C_{n}\}$ for which we predict the summaries $\mathcal{S}=\{S_{1},...,S_{n}\}$. Note, however, that there is an additional translation step involved simplifying the context into layperson language (Weng et al., 2019). On average, the context $C$ was 552.0 and the summary $S$ was 113.2 words long (see Table 4).
We included three models for patient summarization. Further details on parameter and prompt tuning are given in Appendix C and D:
  • LED: The Longformer Encoder-Decoder was used as a baseline model (Beltagy et al., 2020). [^5] Huggingface models allenai/led-{base/large}-16384 The LED model was initialized by BART (Lewis et al., 2020) and can process 16K tokens. BART can only handle 1K tokens, which was insufficient. For training, we used full fine-tuning on all MIMIC-IV-Note-DI examples. The datasets with 100 derived examples were to small.
  • Llama: Llama 2 (Touvron et al., 2023) has shown promising performance on clinical text summarization (Van Veen et al., 2023b), and we used the versions with 7B and 70B parameters. [^6] Huggingface models meta-llama/Llama-2-{7,70}b-hf We always used 100 training examples for parameter-efficient fine-tuning with LoRA (Hu et al., 2021) and loaded the models in 8-bit to reduce the memory usage. Due to training time constraints, we did not train with all examples.
  • GPT-4: GPT-4 (OpenAI et al., 2023) represents the state of the art in clinical summarization (Van Veen et al., 2023b). We accessed the model via the Azure OpenAI service with opt-out for human review of the data to ensure data privacy. We tested the model with 5 in-context examples (5-shot) or no examples (0-shot).

### v2

@@ Abstract
Patients often face difficulties in understanding their hospitalizations, while healthcare workers have limited resources to provide explanations. In this work, we investigate the potential of large language models to generate patient summaries based on doctors’ notes and study the effect of training data on the faithfulness and quality of the generated summaries. To this end, we release (i) a rigorous labeling protocol for errors in medical texts and (ii) a publicly available dataset of annotated hallucinations in 100 doctor-written and 100 generated summaries. We show that fine-tuning on hallucination-free data effectively reduces hallucinations from 2.60 to 1.55 per summary for Llama 2, while preserving relevant information. We observe a similar effect on GPT-4 (0.70 to 0.40), when the few-shot examples are hallucination-free. We also conduct a qualitative evaluation using hallucination-free and improved training data. We find that common quantitative metrics do not correlate well with faithfulness and quality. Finally, we test GPT-4 for automatic hallucination detection, which clearly outperforms common baselines.
@@
We use MIMIC-IV-Note (Johnson et al., 2023; Goldberger et al., 2000). We create a dataset of discharge instructions (MIMIC-IV-Note-Ext-DI) and hallucination annotations for 100 doctor-written and 100 generated patient summaries (Hallucinations-{MIMIC,Generated}-DI). Our data and code: https://doi.org/10.13026/m6hf-dq94 and https://github.com/stefanhgm/patient_summaries_with_llms.
Institutional Review Board (IRB)
Our work did not require IRB approval. The two clinical annotators are authors of this paper, and both had credentialed access to the MIMIC-IV-Note dataset.
@@ 1 Introduction
Figure 1: We developed a protocol for annotating hallucinations in medical text. Following this protocol, two medical experts labeled hallucinations in 100 doctor-written (Hallucinations-MIMIC-DI) and 100 LLM-generated patient summaries (Hallucinations-Generated-DI). We used the labeled hallucinations in the doctor-written summaries to derive two additional datasets by replacing or removing hallucinations (Cleaned) and by further improving the language (Cleaned & Improved). We used these two datasets for our data-centric hallucination reduction and qualitative experiments.
@@
Many patients do not understand the events that occurred during their hospitalization and the subsequent actions they need to take (Kebede et al., 2014). For instance, Horwitz et al. (2013) performed post-discharge interviews and found that only 59.6% of the patients were able to accurately describe their admission diagnosis and 43.9% could fully describe their scheduled follow-up appointments. Improved discharge communication is associated with lower hospital readmission rates and higher adherence to treatment regiments (Becker et al., 2021). A potential intervention to improve patient comprehension could be patient-oriented summaries that describe all relevant facts in layperson language (Federman et al., 2018). However, writing high-quality patient summaries is a difficult and time-consuming task (Mueller et al., 2015), and healthcare workers already face high workloads (Phillips, 2020; Watson et al., 2019).
@@
Large language models (LLMs) have demonstrated strong capabilities on many natural language tasks including medical summarization (Van Veen et al., 2024). However, LLMs are prone to generating unsupported or erroneous facts also referred to as hallucinations (Maynez et al., 2020). In healthcare, this issue is further aggravated by the fragmented nature of healthcare data, as datasets often do not perfectly mimic the data available at the point of care. For example, radiologists often compare to previous images in their reports, even though many medical imaging datasets do not have the historical images available. Similarly, datasets for medical summarization may not include the full patient history to accompany the written summarization. As a result, there may be references in the summary to information that is not supported by the available context or patient history. Training or fine-tuning on this data replicates these artifacts leading to “hallucinations”. Several techniques for preventing hallucinations have been studied (Huang et al., 2023). However, hallucinations can vary highly in complexity, escaping automatic detection and making careful human annotation necessary (Thomson et al., 2023; Moramarco et al., 2022).
@@
Recent work has emphasized the sample efficiency of LLMs. Zhou et al. (2023) showed that 1,000 carefully selected fine-tuning examples sufficed for the successful alignment of Llama 2 (Touvron et al., 2023). An additional set of 30 examples was found to enable multi-turn dialogues (Zhou et al., 2023). Also, GPT-4 achieved state-of-the-art performance on many benchmarks using at most 25 in-context examples (OpenAI et al., 2023). These findings open up opportunities for data-centric approaches that leverage data curated by human domain experts. In this study, we manually create 100 training examples without hallucinations and 100 training examples with improved quality to study their effect on the faithfulness and quality of LLMs. Our main contributions are:
  1. 1.
    We introduce a rigorous annotation protocol for token-level errors in medical texts.
  2. 2.
    We release two datasets of 100 doctor-written (MIMIC) and 100 generated patient summaries with hallucinations labeled by two trained medical experts requiring 70 hours per annotator.
  3. 3.
@@
    For patient summaries, we demonstrate that fine-tuning on data with manually removed hallucinations can effectively reduce the hallucinations in LLMs while preserving key information.
  4. 4.
    We evaluate GPT-4 for automatic hallucination detection using our datasets, emphasizing its suitability for identifying unsupported evidence.
@@ 2 Related Work
Given the large burden of and repetition within clinical documentation, automated clinical summarization is of great interest across healthcare; generation targets include discharge summaries, radiology impression sections, patient-facing instructions, and problem lists (Van Veen et al., 2024; Adams et al., 2022). Consequently, there has been significant research in the past few years exploring how the advent of LLMs can be leveraged to automate clinical summarization. Approaches for control have included both zero- and few-shot prompting (Van Veen et al., 2024) to fine-tuning (Adams et al., 2022, 2023; Moramarco et al., 2022; Cai et al., 2022; Van Veen et al., 2023). Results have been promising; for example, Van Veen et al. (2024) found that GPT-4 is often preferred over human-generated summaries in terms of correctness and completeness.
@@
However, hallucinations – the generation of facts not grounded in the original text — are a recurring concern in text summarization, both across natural language processing (NLP) in general and specifically in the clinical domain (Zhang et al., 2020b; Kryscinski et al., 2020; Adams et al., 2022; Cai et al., 2022; Xie et al., 2023). Given the ubiquity of this issue, several taxonomies have sprung up in the wider NLP community for producing fine-grained annotations of the accuracy of generated text (Thomson and Reiter, 2020; Mishra et al., 2024). In the clinical domain, there have been similar efforts to annotate hallucinations in generated text, but this work has often occurred on proprietary data, with coarse buckets, or without the release of annotations (Adams et al., 2023; Cai et al., 2022; Van Veen et al., 2024). The closest public release was a set of edits clinicians made to summaries auto-generated from transcripts, but this did not target hallucinations specifically (Moramarco et al., 2022). To our knowledge, this work is the first to release a dataset highlighting hallucinations in clinical summarization. These are particularly important because automated metrics of summarization quality, while fast to compute, often do not correlate with manually evaluated performance, a finding also replicated in this work (Van Veen et al., 2024).
@@
There have been three major strategies to mitigate the effect of hallucinations in auto-generated summaries. The first is to conduct post-hoc detection of hallucinations. This has largely involved learning a model of hallucinations from synthetically generated data (Cai et al., 2022; Zhou et al., 2021; Kryscinski et al., 2020) or identifying normalized concepts (e.g., UMLS concepts) that appear in the summary, but not the source (Nan et al., 2021; Adams et al., 2022). The second strategy is to minimize hallucinations by changing the underlying generation model, e.g., by grounding in Wikipedia, or first converting the source context as triples as an intermediary before generation (Zhang et al., 2020b; Tian et al., 2023; Semnani et al., 2023; Aralikatte et al., 2021; Cao et al., 2018; Cao and Wang, 2021). The final strategy is to improve the quality of the training data itself, either by removing poor training examples (Nan et al., 2021) or by improving the quality of the training examples. For example, to avoid decreasing sample size, Adams et al. (2022) revised the reference text via contrastive learning on synthetic data. In this work, we show that for sample-efficient LLMs, it is sufficient to revise only a small number of training examples that are feasible even manually.
@@ 3 Datasets
3.1 MIMIC Discharge Instructions Datasets
Dataset Size Description
MIMIC Discharge Instructions Datasets
MIMIC-IV-Note-Ext-DI 100,175 Summarization dataset derived from MIMIC-IV-Note with the section Discharge Instructions as summary and all prior notes as context
MIMIC-IV-Note-Ext-DI-BHC 100,175 MIMIC-IV-Note-Ext-DI with the Brief Hospital Course as context
MIMIC-IV-Note-Ext-DI-BHC-Anno 26,178 Subset of MIMIC-IV-Note-Ext-DI-BHC with contexts $\leq$ 4,000 characters and summaries $\geq$ 600 characters to facilitate human annotation
Hallucination Datasets Annotated by Two Medical Experts
Hallucinations-MIMIC-DI 100 Random examples from MIMIC-IV-Note-Ext-DI-BHC-Anno
Hallucinations-Generated-DI 100 20 random contexts from M.-IV-Note-Ext-DI-BHC-Anno and summaries generated with five models during hallucination-reduction experiments
Derived Datasets from Hallucinations-MIMIC-DI
Original 100 Context-summary pairs from Hallucinations-MIMIC-DI
@@ 3.1 MIMIC Discharge Instructions Datasets
Cleaned 100 Original with labeled hallucinations manually removed or replaced
Cleaned & Improved 100 Cleaned with mistakes and artifacts removed or corrected
Table 1: Overview of all datasets used in this work. All datasets are publicly available on PhysioNet.
For our experiments, we created a summarization dataset with clinical notes as context and patient summaries written by doctors as targets. We used the MIMIC-IV-Note v2.2 database, which includes 331,793 deidentified clinical notes from 145,915 patients admitted to Beth Israel Deaconess Medical Center in Boston, MA, USA. Each note consists of various sections that describe a patient’s hospital course. We selected the Discharge Instructions (DI) as summaries and the Brief Hospital Course (BHC) as contexts since the BHCs contain the most relevant information for medical professionals (see Figure 2). We chose this shorter context to reduce the effort for the human annotators and to better fit it into the models’ context windows. Many DIs contained irrelevant artifacts that could distort the downstream analysis. For instance, they consisted of static templates or started with a personal salutation. Hence, we designed a preprocessing pipeline selecting and cleaning 100,175 of the original 331,793 MIMIC-IV-Note examples (see Appendix A). The resulting dataset is named MIMIC-IV-Note-Ext-DI-BHC (see Table 1). We also release a version of the dataset with a longer context using all note sections before the DI, including the BHC (MIMIC-IV-Note-Ext-DI). Details for both summarization datasets can be found in Table 6 in the Appendix. To further facilitate human annotation, we considered a subset of the data with context lengths of at most 4,000 characters and summary lengths of at least 600 characters. This was done to reduce the amount of context to take into account for the annotators and to increase the information in the summaries. The resulting subset contained 26,178 entries (MIMIC-IV-Note-Ext-DI-BHC-Anno) and we used it to sample examples for human annotation.
3.2 Hallucination Datasets
We collected hallucination labels for 100 doctor-written and 100 LLM-generated patient summaries. The doctor-written summaries were selected at random from MIMIC-IV-Note-Ext-DI-BHC-Anno. For the LLM-generated summaries, we selected 20 held-out contexts from MIMIC-IV-Note-Ext-DI-BHC-Anno and used five models, which had been trained for the data-centric hallucination reduction experiments, to generate the summaries. Two medical experts labeled hallucinations in both datasets, following our annotation protocol. This resulted in two datasets: Hallucinations-MIMIC-DI for the doctor-written summaries and Hallucinations-Generated-DI for the LLM-generated summaries (see Table 1).
3.3 Derived Datasets
We define Original as a synonym for the data in Hallucinations-MIMIC-DI because it contains the original 100 doctor-written summaries that may include unsupported facts or errors, referred to as hallucinations. Based on the human labels of these hallucinations, we derived two additional datasets from Hallucinations-MIMIC-DI (see Table 1). The Cleaned dataset contains the same patient summaries with annotated hallucinations manually replaced or removed. We used the Original and Cleaned datasets to test the data-centric hallucination reduction approach (see Section 4.4). For the Cleaned & Improved data, we further manually corrected mistakes and artifacts in the summaries. This dataset was used to fine-tune or prompt the models for our qualitative evaluation (see Section 4.6).
@@ 4.1 Generation of Patient Summaries
In this work, we generated patient summaries given the BHC as context. Formally, we have a set of contexts $\mathcal{C}=\{C_{1},...,C_{n}\}$ for which we predict the summaries $\mathcal{S}=\{S_{1},...,S_{n}\}$. Note, however, that there is an additional translation step that simplifies the context into layperson language (Weng et al., 2019). On average, the context $C$ was 552.0 words long, and the summary $S$ was 113.2 words long (see Table 6 in the Appendix).
We included Llama 2 and GPT-4 in the data-centric hallucination reduction experiments since they are commonly used models that allow for sample-efficient alignment (Zhou et al., 2023; OpenAI et al., 2023). Since this experiment required expensive manual annotations of hallucinations, we could not test additional models. For the quantitative and qualitative evaluation, we also included the Longformer Encoder-Decoder (LED), which has shown good performance in medical summarization (Cai et al., 2022; Adams et al., 2022). Further details on parameter tuning for LED and Llama 2, and prompt tuning for GPT-4, are given in Appendices C and D.
  • LED: The Longformer Encoder-Decoder was used as a baseline model (Beltagy et al., 2020). [^1] Huggingface models allenai/led-{base/large}-16384 The LED model was initialized by BART (Lewis et al., 2020) and can process 16K tokens. BART can only handle 1K tokens, which was insufficient. For training, we used full fine-tuning on 80% of MIMIC-IV-Note-Ext-DI-BHC(-Anno). The datasets with 100 examples were to small.
  • Llama: Llama 2 (Touvron et al., 2023) has shown promising performance on clinical text summarization (Van Veen et al., 2024), and we used the versions with 7B and 70B parameters. [^2] Huggingface models meta-llama/Llama-2-{7,70}b-hf We always used 100 training examples for parameter-efficient fine-tuning with LoRA (Hu et al., 2021) and loaded the models in 8-bit to reduce the memory usage. Due to training time constraints, we did not train with all examples.
  • GPT-4: GPT-4 (OpenAI et al., 2023) represents the state of the art in clinical summarization (Van Veen et al., 2024). We accessed the model via the Azure OpenAI service with opt-out for human review of the data to ensure data privacy. We tested the model with 5 in-context examples (5-shot) or no examples (0-shot).
4.2 Protocol for Labeling Hallucinations
We developed a protocol for labeling token-level errors in medical texts based on (Thomson and Reiter, 2020, 2021). [^3] See the Word document on GitHub: https://github.com/ehudreiter/accuracySharedTask/blob/main/example_exercise/Example_Annotation_Exercise.docx Our main focus was to annotate hallucinations in patient summaries. However, we believe that with slight modifications, it could be applicable to other medical scenarios. We distinguished between unsupported, contradicted, and incorrect facts. Since most hallucinations were unsupported facts, we further distinguished them into nine subcategories (see Figure 2). Thomson and Reiter (2020) considered facts from all sources as given; that is, annotators could also use information from the internet to check facts in the summary. In contrast to that, we treated the context (BHC) as the only ground truth about the patient. We chose this approach to reduce the labeling burden, as annotators could not be expected to review all notes and structured information of a patient. However, since patient summaries contain not only patient-specific information, we did allow general medical knowledge and advice even if not explicitly provided in the context (e.g., “Please take your medications as prescribed”). Clarifications that arose during the annotation process were added to the protocol. The final labeling protocol can be found in Appendix LABEL:apd:protocol.